OpenStudy (anonymous):

what is the solution of the equation 6-2^x=1 answer in decimal form

OpenStudy (anonymous):

Subtract both sides by \(6\)

OpenStudy (anonymous):

\[ -2^x=-5 \]

OpenStudy (anonymous):

-2^x = -5 2^x = 5 xlog2 = 5 x = 5 / log 2